OpenStudy (anonymous):

f(k(x)) if k(x)=x^-2/3 and f(x)=2x^3

OpenStudy (anonymous):

3xk(x)=1/x

OpenStudy (campbell_st):

substitute \[x^{-2/3}\] for x i f(x) \[f(x^{-2/3}) = 2(x^{-2/3})^3\] using index laws.... multiply the powers \[f(k(x)) = 2x^{-2}\] or \[f(k(x)) = 2/x^2\]